A Bibliography of Publications of Ramon E. Moore

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 0.30

Abstract

This bibliography records publications of Ramon E. Moore

Title word cross-reference

n [26].

1604 [17]. 1980 [70].

2005 [75]. 23rd [71].

3600 [17].

6-7 [50].

91m [50]. accuracy [37, 46, 50, 52]. ACM [75]. Advances [53, 74]. aided [73, 47, 48]. Airborne [32]. AIRDOS [32]. AIRDOS-EPA [32]. alert [45]. Analysis [2, 6, 33, 60, 69, 73, 1, 11, 15, 19, 42, 44, 47, 48, 57, 61, 62]. Angewandte [70]. Appl. [50]. Applications [29, 33, 16, 9, 14]. Applied [75, 75]. APPN [53]. approximate [37]. Approximation [3]. arbitrary [23]. architecture [53]. Arithmetic [7, 17, 6, 13, 16, 9, 10, 35, 14]. aspects [18]. Automatic [1, 6, 12, 11].

based [11]. bis [68]. Book [62, 63, 3, 5]. Bounded [26]. Bounding [29]. bounds [12]. Br [70].

Carlo [42]. CDC [17]. Chebyshev [41]. Chicago [65]. close [51]. Complex [9, 31]. Comput. [50]. Computation [22, 66, 67, 69, 1, 12, 18, 24, 63]. Computational [44, 60, 21, 30]. Computer [73, 69, 47, 48]. computer-oriented [69]. Computerized [32]. computers [19, 14]. Computing [38, 6, 26, 72, 58, 75, 11, 25, 59]. Concentrations [32]. Conference [65, 71]. Control [71, 11]. convergence [30]. coordinate [12].

Daniel [63]. database [31]. dawning [54]. December [71]. Decision [71]. Difeq [17, 8]. Differential [17, 22, 63, 5, 24, 43, 12, 52]. Digital [6, 66, 67, 1, 11, 14]. Discrete [5]. distributed [31]. Dose [32].

Early [55]. ed [3]. Electronics [65]. elements [25]. engineering [59]. Environmental [32]. EPA [32]. Equations [17, 22, 29, 43, 12, 23, 39, 41, 52, 21, 24, 63, 5]. Erratum [50]. Error [6, 66, 67, 1, 11, 12]. Estimating [32]. existence [28].

Fe [75]. fixed [56]. form [56]. Forschungsinstitut [68]. Freiburg [70]. Function [26, 29]. Functional [19, 60, 44]. Fundamentals [69]. Funktionalanalytische [68]. fuzzy [57].

generalization [39]. Germany [70]. Global [46, 50, 74, 49]. growth [12].

Handling [31]. held [70]. Henrici [5]. Hilton [71]. honor [64].

Illinois [65]. im [68]. Institut [70]. integral [39]. Integrals [4]. integration [8]. International [70]. Interval [7, 17, 10, 2, 4, 6, 15, 33, 34, 40, 43, 72, 47, 48, 57, 70, 13, 16, 14, 9, 11, 12, 18, 35, 41, 61, 70, 62]. Intervallanalyse [20]. Introduction [61]. Introductory [59]. issue [64]. iterative [27, 30].

James [63].

lags [23]. Langer [3]. linear [23, 37]. local [12]. lower [39].

Man [32]. management [45]. manual [8]. March [75]. Math. [50]. Mathematical [25]. mathematics [70]. Mathematik [70, 68]. Mathematischen [68]. Mathematisches [68]. May [70]. McCormick [65]. method [39, 41]. Methoden [68]. Methodology [32]. Methods [5, 33, 34, 43, 72, 27, 30, 42, 49]. Mexico [75]. Microprogrammed [35]. minima [51]. Monte [42]. Moore [62, 63, 55, 64]. MR1096131 [50]. MRC [13]. multivendor [45].

National [65]. networks [45]. Nevada [71]. no. [50]. Nonlinear [29, 34, 40, 28, 30, 36, 37, 41, 21]. Nov [68]. numbers [11]. Numerical [3, 52, 69]. Numerischen [68].

Oberwolfach [68, 68]. October [65]. Operational [7]. Operator [29, 41, 21]. optimization [74, 46, 49, 50]. Order [58]. Ordinary [22, 63, 5, 24, 12]. oriented [69].

Package [13]. Peter [5]. Place [65]. point [56]. Practical [18]. Preface [64]. prescribed [46, 50, 52]. Proceedings [3, 75, 70, 65, 71]. Program [17, 13]. Programming [16]. programs [10]. proofs [73, 47, 48].

queries [31].

R [3]. Radionuclides [32]. Ramon [62, 63]. Range [38, 26]. Rational [26]. Ray [55, 64]. recurrence [23]. reduce [12]. references [10]. Region [26]. regions [27]. Relations [58]. Releases [32]. Reliability [72]. reliable [59]. remarks [59]. resolution [51]. results [36]. Review [62, 63, 3, 5]. Reviews [24]. Rigor [58]. Rigorous [49]. Risk [42]. Role [72]. routine [8].

Safe [27]. Santa [75]. Scientific [72, 25]. set [57]. Sets [29]. simple [37]. SNA [45]. Solution [17, 52, 21]. solutions [12, 28, 37, 39]. some [9, 14]. Spaces [29]. Sparse [56]. Special [64]. stability [23]. starting [27]. Successive [40]. Survey [43]. Symposium [75, 3, 70]. Systems [34, 40, 28, 30, 36, 37, 56].

Tagung [68, 68]. Test [40, 28, 30, 37]. Theory [22, 63, 24, 57]. time [23]. tools [47, 48]. transformations [12].

Universität [70]. upper [39]. USA [75]. use [11]. user [8]. Using [17]. Utilizing [45].

Values [38]. Variable [5]. Variables [26]. Vegas [71]. version [41]. vom [68]. Vortragsauszüge [68].

W [63]. without [42]. works [55].

References

Moore:1959:AEA

[1] R. E. Moore. Automatic error analysis in digital computation. Technical Report Space Div. Report LMSD84821, Lockheed Missiles and Space

Co., Sunnyvale, CA, USA, 1959. URL http://interval.louisiana.edu/
Moores_early_papers/Moore_Lockheed.pdf.

Moore:1959:IAI

[2] R. E. Moore and C. T. Yang. Interval analysis I. Technical Document LMSD-285875, Lockheed Missiles and Space Division, Sunnyvale,

CA, USA, 1959. URL http://interval.louisiana.edu/Moores_early_
papers/Moore_Yang.pdf.

Moore:1960:BRB

[3] Ramon E. Moore. Book review: On Numerical Approximation (Proceedings of a Symposium) (R. E. Langer, ed.). SIAM Review, 2(1):49–50, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1960:II

[4] R. E. Moore, W. Strother, and C. T. Yang. Interval integrals. Technical Memorandum: Mathematics LMSD-703073, Lockheed Missiles and Space Division, Sunnyvale, CA, USA, 1960. URL http://interval.louisiana.

edu/Moores_early_papers/Moore_integrals.pdf.

Moore:1962:BRD

[5] Ramon E. Moore. Book review: Discrete Variable Methods in Ordinary Differential Equations by Peter Henrici. Science, 136(3511):143– 144, April 13, 1962. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://science.sciencemag.org/content/136/
3511/143.2;   http://www.jstor.org/stable/1708446.

Moore:1962:IAA

[6] R. E. Moore. Interval Arithmetic and Automatic Error Analysis in Digital Computing. Ph.D. dissertation, Department of Mathematics, Stanford University, Stanford, CA, USA, November 1962. URL http://interval.

louisiana.edu/Moores_early_papers/disert.pdf. Also published as

Applied Mathematics and Statistics Laboratories Technical Report No. 25.

Boche:1963:OIA

[7] R. E. Boche. An operational interval arithmetic. In IEEE [65],

page ?? URL http://interval.louisiana.edu/Moores_early_papers/
Boche_operational.pdf.

Moore:1964:DIR

[8] R. E. Moore, J. A. Davison, H. R. Jaschke, and S. Shayer. DIFEQ integration routine — user’s manual. Technical Report LMSC6-90-64-6, Lockheed Missiles and Space Co., Los Angeles, CA, 1964. URL http://interval.

louisiana.edu/Moores_early_papers/Moore_DIFEQ.pdf.

Boche:1965:CIA

[9] R. E. Boche. Complex interval arithmetic with some applications. Technical Report LMSC4-22-66-1, Lockheed Missiles and Space Co., Sunnyvale,

CA, USA, 1965. URL http://interval.louisiana.edu/Moores_early_
papers/Boche_complex.pdf.

Miller:1965:IAP

[10] M. E. Miller. Interval arithmetic programs and references. Memo, Lockheed Missiles and Space Div., Sunnyvale, CA, USA, 1965. URL http:

//interval.louisiana.edu/Moores_early_papers/Miller_programs.
pdf.

Moore:1965:AACa

[11] Ramon E. Moore. The automatic analysis and control of error in digital computing based on the use of interval numbers. In Rall [66], chapter 2,

pages 61–130. URL http://interval.louisiana.edu/Moores_early_

papers/Moore_in_Rall_V1.pdf. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Moore:1965:AACb

[12] Ramon E. Moore. Automatic local coordinate transformations to reduce the growth of error bounds in interval computation of solutions of ordinary differential equations. In Rall [67], chapter 2, pages 103–

140. URL http://interval.louisiana.edu/Moores_early_papers/

Moore_in_Rall_V2.pdf. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Reiter:1965:IAP

[13] A. Reiter. Interval arithmetic package: MRC program 2. Coop organ, code-wisc. math. res. center, University of Wisconsin, Madison, 1965.

URL http://interval.louisiana.edu/Moores_early_papers/Reiter_
package.pdf.

Shayer:1965:IAS

[14] S. Shayer. Interval arithmetic with some applications for digital computers. Technical Report LMSD5-13-65-12, Lockheed Missiles and Space Co., Sun-

nyvale, CA, USA, 1965. URL http://interval.louisiana.edu/Moores_
early_papers/Shayer_applications.pdf.

Moore:1966:IA

[15] Ramon E. Moore. Interval analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1966. xi + 145 pp. LCCN QA297 .M63.

Reiter:1967:PIA

[16] A. Reiter. Programming interval arithmetic and applications. In Proceedings of the 1967 Army Numerical Analysis Conference, pages 87–98. U. S. Army Research Office, Durham, NC, USA, 1967. URL http://interval.

louisiana.edu/Moores_early_papers/Reiter_programming.pdf. Army

Research Office Rep. 67-3.

Braun:1968:PSD

[17] J. A. Braun and R. E. Moore. A program for the solution of differential equations using interval arithmetic (difeq) for the CDC 3600 and the CDC 1604. MRC Technical Summary 901, University of Wisconsin,

Madison, 1968. URL http://interval.louisiana.edu/Moores_early_
papers/addendum/Braun1968aprogram.pdf.

Moore:1968:PAI

[18] R. E. Moore. Practical aspects of interval computation. Aplikace Matematiky, 13(??):52–92, 1968. CODEN APMTAK. ISSN 0373-6725.

Moore:1969:FAC

[19] Ramon E. Moore. Functional analysis for computers. In und Heinz Unger [68], pages 113–126. LCCN QA297 .T3 1967aa.

Moore:1969:IUD

[20] Ramon E. Moore. Intervallanalyse. R. Oldenbourg, München, Germany, 1969. DM 39.00. Translated to German by Dieter Pfaffenzeller.

Rall:1969:CSN

[21] Louis B. Rall. Computational solution of nonlinear operator equations. With an appendix by Ramon E. Moore. Wiley, New York, NY, USA, 1969. viii + 225 pp.

Daniel:1970:CTO

[22] James W. Daniel and Ramon E. Moore. Computation and Theory in Ordinary Differential Equations. W. H. Freeman, San Francisco, CA, USA, 1970. ISBN 0-7167-0440-4. xi + 172 pp. LCCN QA372 .D28. Sec. 5.8 (pp. 86–89) and Sec. 6.6 (pp. 100–101).

Moore:1970:SLR

[23] Ramon E. Moore. On the stability of linear recurrence equations with arbitrary time lags. Journal of Computer and System Sciences, 4(4):377– 383, August 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Moore:1972:RCT

[24] Ramon E. Moore, James W. Daniel, and W. E. Boyce. Reviews: Computation and Theory in Ordinary Differential Equations. American Mathematical Monthly, 79(4):407–408, 1972. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Moore:1975:MES

[25] Ramon E. Moore. Mathematical elements of scientific computing. Holt, Rinehart, and Winston, New York, NY, USA, 1975. ISBN 0-03-088125-0. x + 237 pp. LCCN QA 297 .M64 1975.

Moore:1976:CRR

[26] R. E. Moore. On computing the range of a rational function of n variables over a bounded region. Computing: Archiv für informatik und numerik, 16 (1–2):1–15, 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057

(electronic). URL http://springerlink.metapress.com/openurl.asp?
genre=journal&issn=0010-485X.

Moore:1977:SSR

[27] R. E. Moore and S. T. Jones. Safe starting regions for iterative methods. SIAM Journal on Numerical Analysis, 14(6):1051–1065, 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1977:TES

[28] R. E. Moore. A test for existence of solutions to nonlinear systems. SIAM Journal on Numerical Analysis, 14(??):611–615, 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1978:BSF

[29] R. E. Moore. Bounding sets in function spaces with applications to nonlinear operator equations. SIAM Review, 20(3):492–512, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1978:CTC

[30] R. E. Moore. A computational test for convergence of iterative methods for nonlinear systems. SIAM Journal on Numerical Analysis, 15(6):1194–1196, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore79

[31] R. E. Moore. Handling complex queries in a distributed database. Technical Note, Artificial Intelligence 170, SRI International, Menlo Park, CA, USA, 1979. ???? pp.

Moore:1979:AEC

[32] R. E. Moore et al. AIRDOS-EPA: A computerized methodology for estimating environmental concentrations and dose to man from airborne releases of radionuclides. Technical Report ORNL-5532, Oak Ridge National Laboratory, Oak Ridge, Tenn., 1979.

Moore:1979:MAI

[33] Ramon E. Moore. Methods and Applications of Interval Analysis. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-161-4. xi + 190 pp. LCCN QA297.75 .M66.

Moore:1980:IMN

[34] R. E. Moore. Interval methods for nonlinear systems. In Alefeld and Grigorieff [69], pages 113–120. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Moore:1980:MIA

[35] R. E. Moore. Microprogrammed interval arithmetic. ACM SIGNUM Newsletter, 15(2):30, June 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moore:1980:NRN

[36] Ramon E. Moore. New results on nonlinear systems. In Nickel [70], pages 165–180. ISBN 0-12-518850-1. LCCN QA297.75 .I57 1980.

Moore:1980:STA

[37] R. E. Moore and J. B. Kioustelidis. A simple test for accuracy of approximate solutions to nonlinear (or linear) systems. SIAM Journal on Numerical Analysis, 17(4):521–529, August 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Asaithambi:1982:CRV

[38] N. S. Asaithambi, Shen Zuhe, and R. E. Moore. On computing the range of values. Computing: Archiv für informatik und numerik, 28(3): 225–237, 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057

(electronic). URL http://springerlink.metapress.com/openurl.asp?
genre=journal&issn=0010-485X.

Moore:1982:GMU

[39] R. E. Moore. A generalization of the method of upper and lower solutions for integral equations. Nonlinear Anal., Theory Methods Appl., 6(??):829– 831, 1982.

Moore:1982:SIT

[40] R. E. Moore and L. Qi. A successive interval test for nonlinear systems. SIAM Journal on Numerical Analysis, 19(4):845–850, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1983:IVC

[41] R. E. Moore and Zuhe Shen. An interval version of chebyshev’s method for nonlinear operator equations. Nonlinear Anal., Theory Methods Appl., 7(??):21–34, 1983.

Moore:1984:RAM

[42] Ramon E. Moore. Risk analysis without Monte Carlo methods. Freiburger Intervall-Berichte 84/1, Universität Freiburg, Freiburg, Germany, 1984. 48 pp.

Moore:1984:SIM

[43] Ramon E. Moore. A survey of interval methods for differential equations. In IEEE [71], pages 1529–1535. ISBN ???? LCCN TJ217 .I17 1984. Three volumes.

Moore:1985:CFA

[44] Ramon E. Moore. Computational functional analysis. Mathematics and its Applications. Ellis Horwood, New York, NY, USA, 1985. ISBN 0-85312-807-3. 156 pp.

Moore:1988:USA

[45] R. E. Moore. Utilizing the SNA alert in the management of multivendor networks. IBM Systems Journal, 27(1):15–31, January 1988. CODEN IBMSA7. ISSN 0018-8670.

Moore:1991:GOP

[46] Ramon E. Moore. Global optimization to prescribed accuracy. Computers and Mathematics with Applications, 21(6–7):25–39, 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Moore:1991:ITC

[47] Ramon E. Moore. Interval tools for computer aided proofs in analysis. In Meyer and Schmidt [73], pages 211–216. ISBN 0-387-97426-1 (New York Berlin Heidelberg alk. paper) 3-540-97426-1. LCCN ???? (Berlin Heidelberg New York: alk. paper).

Moore:1991:TMC

[48] Ramon E. Moore. Interval tools for computer aided proofs in analysis. In Meyer and Schmidt [73], page ?? ISBN 0-387-97426-1 (New York Berlin Heidelberg alk. paper) 3-540-97426-1. LCCN ???? (Berlin Heidelberg New York: alk. paper).

Moore:1992:RMG

[49] Ramon Moore, Eldon Hansen, and Anthony Leclerc. Rigorous methods for global optimization. In Floudas and Pardalos [74], pages 321–342. ISBN 0-691-08740-7 (hardback), 0-691-02527-4 (paperback). LCCN QA402.5 .R42 1992. Papers presented at a conference held at Princeton University, May 10–11, 1991.

Moore:1993:EGO

[50] Ramon E. Moore. Erratum to: “Global optimization to prescribed accuracy” [Comput. Math. Appl. 21 (1991), no. 6-7, 25–39; MR1096131 (91m:90159)]. Computers and Mathematics with Applications, 25(10–11): 187, 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Moore:1993:RCM

[51] R. E. Moore. The resolution of close minima. Computers and Mathematics with Applications, 25(10–11):57–58, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/089812219390281Y.

Moore:1994:NSD

[52] R. E. Moore. Numerical solution of differential equations to prescribed accuracy. Computers and Mathematics with Applications, 28(10–12):253–261, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Bird:1995:AAA

[53] B. R. Bird, C. Brotman, R. Case, G. Dudley, R. E. Moore, and M. Peters. Advances in APPN architecture. IBM Systems Journal, 34(3):430–451, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.research.

ibm.com/journal/sj34-3.html#seven.

Moore:1999:D

[54] R. E. Moore. The dawning. Reliable Computing = Nadezhnye vychisleniia, 5:423–424, 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Hansen:2001:EWR

[55] Eldon Hansen, Bill Walster, Ray Moore, et al. Early works of Ray Moore. World-Wide Web document, September 2001. URL http:/

/interval.louisiana.edu/Moores_early_papers/bibliography.html.

From an announcement by R. Baker Kearfott <rbk@louisiana.edu> to the

reliable computing@interval.louisiana.edu list on Fri, 28 Sep 2001

19:56:51 -0500: “Bill Walster has collected the early works of Ray Moore and his colleagues, has scanned them into electronic form, and has obtained permission from the publishers to post them. (Sun has sponsored this endeavor, including paying copyright fees.) Eldon Hansen has written a short introduction, that I have converted to HTML. This HTML file contains links to the actual papers (in PDF format), that you can access and read.”.

Moore:2002:SSF

[56] Ramon E. Moore. Sparse systems in fixed point form. Reliable Computing = Nadezhnye vychisleniia, 8(4):249–265, 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Moore:2003:IAF

[57] Ramon Moore and Weldon Lodwick. Interval analysis and fuzzy set theory. Fuzzy Sets and Systems, 135(1):5–9, 2003. CODEN FSSYD8. ISSN 0165-0114 (print), 1872-6801 (electronic).

Moore:2005:ORR

[58] Ramon E. Moore. Order relations and rigor in computing. In Haddad et al. [75], pages 1431–1433. ISBN 1-58113-964-0. LCCN QA76.76.A65 S95 2005. Keynote address.

Moore:2006:IRR

[59] Ramon E. Moore. Introductory remarks on reliable engineering computing. Reliable Computing = Nadezhnye vychisleniia, 12(6):405–408, 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Moore:2007:CFA

[60] Ramon E. Moore and Michael J. Cloud. Computational Functional Analysis. Ellis Horwood, New York, NY, USA, second edition, 2007. ISBN 1-904275-24-9. xii + 180 pp. LCCN ???? US$60. URL http://www.

horwoodpublishing.net/bookpage.php?id=110.

Moore:2009:IIA

[61] Ramon E. Moore, R. Baker Kearfott, and Michael J. Cloud. Introduction to interval analysis. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2009. ISBN 0-89871-669-1. xii + 223 pp.

Hansen:1967:BRB

[62] Eldon Hansen. Book review: Interval Analysis (Ramon E. Moore). SIAM Review, 9(3):610–612, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hull:1971:BRB

[63] T. E. Hull. Book review: Computation and Theory in Ordinary Differential Equations (James W. Daniel and Ramon E. Moore). SIAM Review, 13(3): 413–414, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kearfott:2016:PSI

[64] R. Baker Kearfott. Preface [Special issue in honor of Ray Moore, 1929– 2015]. Reliable Computing = Nadezhnye vychisleniia, 23:1, 2016. CODEN RCOMF8. ISSN 1573-1340.

IEEE:1963:PNE

[65] IEEE, editor. Proceedings of the National Electronics Conference (Mc-Cormick Place, Chicago, Illinois October 28–30, 1963). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1963.

Rall:1965:EDCa

[66] L. B. Rall, editor. Error in Digital Computation, volume 1. Wiley, New York, NY, USA, 1965. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Rall:1965:EDCb

[67] L. B. Rall, editor. Error in Digital Computation, volume 2. Wiley, New York, NY, USA, 1965. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Unger:1969:TFM

[68] Lothar Collatz und Heinz Unger, editor. Tagung über Funktionalanalytische Methoden der Numerischen Mathematik (1967: Mathematisches Forschungsinstitut Oberwolfach) Funktionalanalytische Methoden der numerischen Mathematik. Vortragsauszüge der Tagung über funktionalanalytische Methoden der numerischen Mathematik vom 19. bis 25. Nov. 1967 im Mathematischen Forschungsinstitut Oberwolfach. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1969. LCCN QA297 .T3 1967aa.

Alefeld:1980:FNC

[69] G. Alefeld and R. D. Grigorieff, editors. Fundamentals of numerical computation (computer-oriented numerical analysis), volume 2 of Computing. Supplementum. Springer, Wien / New York, 1980. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Nickel:1980:IMP

[70] Karl L. E. Nickel, editor. Interval mathematics 1980: proceedings of an International Symposium on Interval Mathematics, held at the Institut für Angewandte Mathematik, Universität Freiburg i. Br., Germany, May 27– 31, 1980. Academic Press, New York, NY, USA, 1980. ISBN 0-12-518850-1. LCCN QA297.75 .I57 1980.

IEEE:1984:PIC

[71] IEEE, editor. Proceedings of the 23rd IEEE Conference on Decision and Control, December 12–14, 1984, Las Vegas Hilton, Las Vegas, Nevada.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. ISBN ???? LCCN TJ217 .I17 1984. Three volumes.

Moore:1988:RCR

[72] Ramon E. Moore, editor. Reliability in Computing: the Role of Interval Methods in Scientific Computing, volume 19 of Perspectives in computing. Academic Press, New York, NY, USA, 1988. ISBN 0-12-505630-3. xv + 428 pp. LCCN QA76.9.E94 R45 1988.

Meyer:1991:CAP

[73] Kenneth R. (Kenneth Ray) Meyer and Dieter S. Schmidt, editors. Computer aided proofs in analysis, volume 28 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97426-1 (New York Berlin Heidelberg alk. paper) 3-540-97426-1. LCCN ???? (Berlin Heidelberg New York: alk. paper).

Floudas:1992:RAG

[74] Christodoulos A. Floudas and Panos M. Pardalos, editors. Recent advances in global optimization, Princeton series in computer science. Princeton University Press, Princeton, NJ, USA, 1992. ISBN 0-691-08740-7 (hardback), 0-691-02527-4 (paperback). LCCN QA402.5 .R42 1992. Papers presented at a conference held at Princeton University, May 10–11, 1991.

Haddad:2005:ACP

[75] Hisham M. Haddad et al., editors. Applied computing 2005: proceedings of the 2005 ACM Symposium on Applied Computing, Santa Fe, New Mexico, USA, March 13–17, 2005. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-58113-964-0. LCCN QA76.76.A65 S95 2005.